OpenStudy (anonymous):

A ball is dropped from a height of 10 m. It rebounds 1/2 the distance on each bounce. What is the total distance the ball travels? (Round your answer to the nearest whole number)

OpenStudy (anonymous):

this can be seen as a geometric series with a starting term of 10, and a common ratio of 1/2

OpenStudy (anonymous):

S=u1/(1-r) =10/(1-1/2) =10/0.5 =20

OpenStudy (anonymous):

20m using the formula for geometric infinite series

OpenStudy (anonymous):

so the \[Sum _{\infty} = \frac{10}{1 - \frac{1}{2}}\] which is 20